Big Al's Meat Market has annual sales of $546,000 and cost of goods sold of $370,000. The profit margin is 5.8 percent and the accounts payable period is 31.7 days. What is the average accounts payable balance? Assume 365 days per year. $67,697.16 $36,724.85 $32,134.25 $34,429.55 $29,456.39
Answer #1

For finding the average accounts payable we need to use the formula:

Accounts payable period = (Average accounts payable / Cost of goods sold) * 365

31.7 = (Average accounts payable / 370,000) * 365

Average accounts payable / 370,000 = 31.7 / 365

Average accounts payable = (31.7/365) x 370,000 = $32,134.25